Producing one yard of cloth requires 1 worker and 3 acres of land, while producing one pound of food requires 1 worker and 4 acres of land. Both countries have 50 workers: the U.S. has 160 acres of land, while the ROW has 180. The price of food is always $36/pound, the price of cloth is $30/yard in the U.S. in autarky and $33/yard in both countries under free trade.

Q1 Compare the relative abundance of factors across countries.

Q2 Compare the relative intensity of factor use across goods.

Q3 Determine the pattern of comparative advantage and the pattern of trade.
Answer #1

Ans

1 usa is labour abundant since it's initial price of cloth us less than international price. This means ROW Is land abundant

2 land intensity=3/1=3 in cloth and 4/1=4 in food

3 usa has comparative advantage in cloth as it is labour abundant and thus price of cloth is lower in usa. Row has comparative advantage in food. Usa will export cloth and import food. Row will export food and import cloth